OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1, 3) and (1, 11)?

OpenStudy (2bornot2b):

y-3={(-1-1)/(3-11)}(x+1)

OpenStudy (anonymous):

\[ \frac{y1-y2}{x1-x2}=4\] y=mx+b y=4x+b 3=4(-1)+b 3+4=b b=7 y=4x+7